# MSegRNN:Enhanced SegRNN Model with Mamba for Long-Term Time Series Forecasting

### v2

@@ Abstract
The field of long-term time series forecasting demands handling extensive look-back windows and long-range prediction steps, posing significant challenges for RNN-based methodologies. Among these, SegRNN, a robust RNN-driven model, has gained considerable attention in LTSF analysis for achieving state-of-the-art results while maintaining a remarkably streamlined architecture. Concurrently, the Mamba structure has demonstrated its advantages in small to medium-sized models due to its capability for information selection. This study introduces a variant of SegRNN that preprocesses information using a fine-tuned single-layer Mamba structure. Additionally, it incorporates implicit segmentation and residual structures into the model’s encoding section to further reduce the inherent data iterative cycles of RNN architectures and implicitly integrate inter-channel correlations. This variant, named MSegRNN, utilizes the Mamba structure to select useful information, resulting in a transformed sequence. The linear-strategy-adapted derivative retains the superior memory efficiency of the original SegRNN while demonstrating enhanced performance. Empirical evaluations on real-world LTSF datasets demonstrate the superior performance of our model, thereby contributing to the advancement of LTSF methodologies.
@@ 1 Introduction
In the context of long-term time series analysis, it is essential to capture the sequential relationships between data. Commonly adopted feature representation methods in time series analysis include Convolutional Neural Networks (Wang et al., 2022; Wu et al., 2022; Foumani et al., 2021), Multilayer Perceptrons (Das et al., 2023; Chen et al., 2023; Zeng et al., 2023), attention-based encoding layers (Nie et al., 2022; Wu et al., 2021; Zhang and Yan, 2022), and Recurrent Neural Networks networks based on sequence structure (Lin et al., 2023). CNN-based methods use convolutional kernels to capture temporal and spatial sequences (Foumani et al., 2021); however, the limited size of these kernels prevent the model from effectively capturing long-range dependencies in time series, thus limiting the expressiveness of the feature representation. Transformer-based methods, leveraging attention mechanisms to improve the capture of spatiotemporal sequences, thereby enhancing feature representation capabilities. MLPs have been exrensively applied in time series prediction, with models such as DLinear (Zeng et al., 2023), which use simple linear layers and channel-independent methods, achieving superior performance compared to the advanced Transformer architectures of its time, prompting a reevaluation of the need for complex Transformer structures in time series prediction.
@@
RNN-based methods have achieved remarkable success in long-term time series analysis tasks. The inherent sequential architecture of RNNs enables them to handle short-term time series prediction tasks effectively. However, as the length of the predicted sequence increases, the model’s prediction error rapidly escalates with the propagation of features through the RNN recurrent units (Yunpeng et al., 2017). To address this limitation, Shengsheng Lin et al. (Lin et al., 2023) proposed the segmentation-based SegRNN model. This model aims to minimize the number of RNN recurrent iterations while retaining the sequential information, achieving significant success in long-term time series analysis tasks. SegRNN employ a channel-independent strategy to enhance its effectiveness. Nevertheless, this strategy fails to effectively handle noise and capture the correlations between variables. Additionally, even with the adoption of the segmentation strategy, SegRNN still faces the issue of information loss within the recurrent structure.
In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segment strategy. Specifically, we propose a long-term time series prediction model based on Mamba transformation and linear encoding, termed MSegRNN, the main contributions of MSegRNN can be summarized as follows:
  • Introducing the Mamba structure to effectively extract the information
  • Utilizes information between channels to promote the infomation
@@
The structure of this paper is organized as follow: Section 2 formally presents our problem statement. Section 3 discusses related work. Section 4 elaborates on our method of introducing Mamba structure and other technologys. Section 5 describes our experiments and results. Section 6 concludes the paper and suggests future research directions.
@@ 2 Problem statement
In which $X\in\mathbb{R}^{L\times C}$ represents multi-variable time sequences with historical context, $F$ represents the trained neural network structure, $Y\in\mathbb{R}^{H\times C}$ represents the future time sequences to be predicted, $L$ indicates the historical lookback window, $C$ indicates the number of variables, and H indicates the time step to be predicted. Long-term time series analysis aims to expand the potential of predicting time steps, which poses significant challenges for training neural networks.
3 Related Work
A significant amount of work has been dedicated to advancing the development of deep learning in time series analysis. In this section, we introduce literature relevant to our work. These methods can be broadly divided into the following four categories:
@@ 3.1 Transformer Models
Transformer models, originally designed for sequence-to-sequence tasks, have achieved significant achievements accross various domains. The core of applying Transformer to long-time series tasks is to use the attention mechanism to capture the long-term dependencies (Hao and Cao, 2020). Extensive research has been devoted to advancing the application of the Transformer architecture in long-time series analysis. Autoformer (Wu et al., 2021) decomposes the sequence into basic blocks within the model, thus gradually decomposing complex time series and enhancing learning ability for complex patterns. Informer (Zhou et al., 2021) constructs a ProbSparse self-attention mechanism, adopting self-attention distillation and a generative decoder to address the high time complexity, large memory consumption, and encoder architecture limitations faced in long-time series analysis. Crossformer (Zhang and Yan, 2022) embeds time series into two-dimensional vector through the Dimension-segment-wise structure, effectively captures dependencies across time and dimensions through the Two-stage-attention structure. The segmentation technique used in Crossformer method is considered beneficial for long time series foresting.
@@ 3.2 MLP Models
Extensive research has been dedicated to the application of MLP methods in long-term time series forecasting. Dlinear (Zeng et al., 2023) decomposes time series through the application of moving average kernels and remainder components, yielding two feature variables. These variables are then mapped to the prediction results via linear layers. The Nlinear (Zeng et al., 2023) method adopts a regularization strategy in time series forecasting tasks to enhance performance. Specifically, it subtracts the last value before inputting to the model, mapping through a linear layer, then adding it back for the final prediction. The TiDE (Nie et al., 2022) method introduces a residual structure within the linear layer to encode past time series and covariates, subsequently decoding the encoded time series with future covariates to obtain prediction results. Long-term time series analysis methods based on MLP have demonstrated outstanding predictive accuracy while maintaining low time-space complexity, thus advancing the development of MLP in long-term time series forecasting.
@@ 3.3 CNN Models
The core of time series analysis based on CNN architectures lies in utilizing convolutional kernels to extract temporal and channel features. The Disjoint-CNN (Foumani et al., 2021) approach, while maintaining computational efficiency, enhances prediction accuracy by decomposing one-dimensional convolutional kernels into disjoint temporal and spatial components. Nonetheless, CNN-based methodologies encounter challenges as the forecasting horizon expands. MICN (Wang et al., 2022) aims to fully exploit the latent information in time series by designing a multi-scale hybrid decomposition module to decompose the complex patterns of input time series. Each module within this design leverages subsampling convolution and equidistant convolution to respectively extract local characteristics and global correlations. TimesNet (Wu et al., 2022) transforms one-dimensional time series into a set of two-dimensional tensors based on multiple periodicities. It further introduces the TimesBlock to adaptively uncover these multi-periodic patterns and effectively extracts intricate temporal dynamics from the transformed tensors through parameter-efficient initializations. These methodologies have attained remarkable results in the domain of long-term time series forecasting.
3.4 RNN Models
The RNN architecture was initially designed for processing sequence tasks and excel in short-term time series prediction. However, RNN-based methods face error accumulation issues in long-term time series tasks. SegRNN (Lin et al., 2023) aims to mitigate the accumulation of errors in the recurrent structure while preserving sequence information. Specifically, in the encoding stage, the SegRNN method segments the data before encoding. In the decoding stage, it employs a PMF approach for parallel decoding, significantly enhancing prediction accuracy and the spatiotemporal efficiency.
Nonetheless, the SegRNN method is not without its limitations: (1) SegRNN couldn’t effectivly preporcess the data and extract the inforamtion, which may lead to a rapid accumulation of noise. (2)SegRNN adopts fixed segmentation of the original time series, which may result in partial information loss due to its rigidity. (3) SegRNN aims to minimize the number of recurrent iterations while striving to retain sequential information. However, there are still limitations in minimizing the number of recurrent iterations, which restrict the model’s capability. In the following section, we will propose some strategies to address these issues.
@@ 4 Method
To address the challenges faced by current methods, we incorporate mamba structure, implicit segmentation and residual networks into the SegRNN framework. This approach (1) first extracts the original data first to contrain noise, (2)segments the original data through linear mapping to reduce information loss caused by fixed segmentation, and (3) introduces residual structures at the encoding layer to mitigate information loss within the recurrent structure.
We first discuss the overall sturcture that we used in Section 4.1, then we discuss the comparison between explicit segmentation and implicit segmentation in Section 4.2, followed by a discussion on how to introduce residual structures into the model during the encoding process of time series data in Section 4.3.
4.1 Model Structure
The Mamba architecture achieves long-term sequence memory and model selection capabilities by setting the SSM structure parameters as trainable parameters dependent on the sequence content. We have adapted the Mamba architecture to address time series preprocessing problems. The basic structure of our model is shown in Figure 1:
@@ 4.1 Model Structure
Several modifications were applied to the Mamba block and SegRNN model to improve performance and flexibility. Specifically, the convolutional layer in Mamba was removed to enhance computational efficiency, with local information being handled by the SegRNN structure. The SegRNN model’s fixed segmentation was replaced with Implicit Segmentation for adaptability, and residual connections were integrated to strengthen the linkage between segmentation and encoding outputs. Further details on these adjustments are discussed in Section 4.2 and Section 4.3.
@@ 4.2 Implicit Segmentation
The explicit segmentation approach employed by SegRNN pre-processes the data such that a given multivariate time series, denoted as $X^{(i)}\in\mathbb{R}^{L\times C}$, As shown in Figure 2, where $L$ represents the look-back length and $C$ represents the number of channels in multivariate time series. Upon selecting a segment length $w$, the time series can be transformed into segmented time series $X_{w}^{(i)}\in\mathbb{R}^{n\times w\times C}$, where $n$ represents the number of segments:
@@
Subsequently, the segmented model projects through a linear layer:
$\tilde{X}_{w}^{(i)}=W_{prj}\cdot X_{w}^{(i)}$ (3)
where $W_{prj}\in\mathbb{R}^{w\times d}$ denotes the projection matrix of linear mapping, and $d$ represents the dimensionality of the hidden layer within the RNN structure. Undergoing linear projection transformation, the dimensionality of $\tilde{X}_{w}^{(i)}$ is defined as $\mathbb{R}^{C\times n\times d}$.
However, this segmentation strategy has certain limitations. Once the segment length is chosen, the cut-off points are also determined. This may disrupt the continuity of information in time series analysis, thus limiting the model’s capabilities. The implicit segmentation method aims to break this limitation. Specifically, given a time series analysis vector $X^{(i)}\in\mathbb{R}^{L\times C}$, we first expand its last dimension and then swap the first and second dimensions to obtain $\tilde{X}^{(i)}\in\mathbb{R}^{C\times L\times 1}$ for subsequent processing. Then we utlize a linear layer to transform the last dimesntion of $\tilde{X}^{(i)}$ to get $\bar{X}(i)\in\mathbb{R}^{C\times L\times n}$ to represent the segmented matrix, subsequently, we interchange the second and third dimension, still denoted as $\bar{X}(i)$. Applying linear layer to the last dimension of $\bar{X}(i)$ yields $\bar{X}(i)\in\mathbb{R}^{C\times n\times d}$, which is consistent in dimension with the vector obtained via SegRNN methodology.
Figure 2: Comparsion of SegRNN Segmentation with Implicit Segmentation
Notably, our proposed method does not explicitly define the segmentation length. Instead, it achieves data segmentation and dimensionality transformation through dual linear projections. The incorporation of linear fully connected layers facilitates a more continuous and dense transformation of information. It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe approach but rather represents a general and straightforward technique for information enhancement.
4.3 Residual Structure for SegRNN
SegRNN aims to minimize the number of iterations in the RNN while preserving sequential information. Extensive experiments show that increasing the number of iterations negatively impacts information retention. SegRNN uses data segmentation to reduce the data’s flow through the RNN structure. Specifically, for the vector obtained through segmentation in SegRNN, $\tilde{X}_{w}^{(i)}\in\mathbb{R}^{C\times n\times d}$, we transform it in the RNN structure with a hidden layer dimension $d$ using the following formula:
@@ 4.3 Residual Structure for SegRNN
where $\mathbf{h}_{t}\in\mathbb{R}^{1\times C\times d}$ represents the hidden state at step $t$, $\tilde{\mathbf{X}}_{w}^{(i)}[:,t,:]$ represents the input vector at step $t$, $\mathbf{W}$ denotes the RNN’s weight parameters, and $f$ is the RNN’s activation function.
Using segmentation techniques, SegRNN reduces the number of iterations from $L$ to $n$. However, the number of iterations still increases with the forecasting time steps. This unavoidably limits the model’s ability due to the loss of information during iterations. Therefore, we introduce a residual structure in the model, as shown in Figure 3. For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbb{R}^{C\times L\times n}$, we introduce a linear mapping from $\bar{\mathbf{X}}^{(i)}$ to $\mathbf{h}_{n}$. Specifically, we first merge the last two dimensions and extend the first dimension, obtaining the three-dimensional matrix $\mathbf{X}^{\prime(i)}\in\mathbb{R}^{1\times C\times(L\times n)}$. Then, we transform the last dimension through a linear layer to get $\mathbf{h}^{\prime}_{n}\in\mathbb{R}^{1\times C\times d}$. Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$, obtaining the final output vector from the encoder.
Figure 3: Residual block for SegRNN
Introducing the residual structure allows some information to bypass the recurrent structure and be directly mapped to the encoder output. This approach minimizes information loss within the recurrent structure, providing a beneficial enhancement for time series data.

### v3

@@ Abstract
Long time series forecasting aims to utilize historical information to forecast future states over extended horizons. Traditional RNN-based series forecasting methods struggle to effectively address long-term dependencies and gradient issues in long time series problems. Recently, SegRNN has emerged as a leading RNN-based model tailored for long-term series forecasting, demonstrating state-of-the-art performance while maintaining a streamlined architecture through innovative segmentation and parallel decoding techniques. Nevertheless, SegRNN has several limitations: its fixed segmentation disrupts data continuity and fails to effectively leverage information across different segments, the segmentation strategy employed by SegRNN does not fundamentally address the issue of information loss within the recurrent structure. To address these issues, we propose the MSegRNN method with three key enhancements: we introduce an implicit segmentation structure to decompose the time series and map it to segmented hidden states, resulting in denser information exchange during the segmentation phase. Additionally, we incorporate residual structures in the encoding layer to mitigate information loss within the recurrent structure. To extract information more effectively, we further integrate the Mamba architecture to enhance time series information extraction. Experiments on several real-world long time series forecasting datasets demonstrate that our model surpasses the performance of current state-of-the-art models.
@@ 1 Introduction
In the context of long-term time series analysis, it is essential to capture the sequential relationships between data. Commonly adopted feature representation methods in time series analysis include Convolutional Neural Networks (Wang et al., 2022; Wu et al., 2022; Foumani et al., 2021), Multilayer Perceptrons (Das et al., 2023a; Chen et al., 2023; Zeng et al., 2023), attention-based encoding layers (Das et al., 2023b; Nie et al., 2022; Zhang and Yan, 2022), and Recurrent Neural Networks networks based on sequence structure (Lin et al., 2023). CNN-based methods use convolutional kernels to capture temporal and spatial sequences (Foumani et al., 2021); however, the limited size of these kernels prevent the model from effectively capturing long-range dependencies in time series, thus limiting the expressiveness of the feature representation. Transformer-based methods, leveraging attention mechanisms to improve the capture of spatiotemporal sequences, thereby enhancing feature representation capabilities. MLPs have been exrensively applied in time series prediction, with models such as DLinear (Zeng et al., 2023), which use simple linear layers and channel-independent methods, achieving superior performance compared to the advanced Transformer architectures of its time, prompting a reevaluation of the need for complex Transformer structures in time series prediction.
@@
RNN-based methods have achieved remarkable success in long-term time series analysis tasks. The inherent sequential architecture of RNNs enables them to handle short-term time series prediction tasks effectively. However, as the length of the predicted sequence increases, the model’s prediction error rapidly escalates with the propagation of features through the RNN recurrent units (Yunpeng et al., 2017). To address this limitation, Shengsheng Lin et al. (Lin et al., 2023) proposed the segmentation-based SegRNN model. This model aims to minimize the number of RNN recurrent iterations while retaining the sequential information, achieving significant success in long-term time series foreasting tasks. SegRNN employs a segmented approach to time series, which may disrupt the continuity of the sequences. Although the segmentation strategy effectively balances sequence information and iteration count, information loss within the recurrent structure occurs as the prediction horizon extends, thereby limiting the model’s capability.
In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segmentation strategy. Specifically, we propose a long-term time series forecasting model that utilizes the Mamba structure, implicit segmentation, and residual blocks, termed MSegRNN, the main contributions of MSegRNN can be summarized as follows:
  • Introducing the Mamba structure to effectively preprocessing the information.
  • Utilizing implicit segmentation for denser information exchange during the segmentation phase.
@@ 2 Problem Statement
In which $X\in\mathbf{R}^{L\times C}$ represents multivariable time sequences with historical context, $F$ represents the trained neural network structure, $Y\in\mathbf{R}^{H\times C}$ represents the future time sequences to be predicted, $L$ indicates the historical lookback window, $C$ indicates the number of variables, and $H$ indicates the time step to be predicted. Long-term time series analysis aims to expand the potential of predicting time steps, which poses significant challenges for training neural networks.
A significant amount of work has been dedicated to advancing the development of deep learning in time series analysis. These methods can be broadly divided into the following categories:
@@ 2.1 Transformer Models
The core of applying Transformer models to long-term time series tasks lies in their ability to capture long-term dependencies using the attention mechanism (Hao and Cao, 2020). The general form of the self-attention mechanism is given by:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (2)
Significant research efforts have advanced the application of Transformer architectures in long-term time series analysis. For instance, PatchTST(Nie et al., 2022) introduces a method of segmenting time series into patches, reducing computational complexity and improving the model’s ability to capture local semantic information. It also employs self-supervised learning, masking random patches and training the model to reconstruct them, enhancing forecasting accuracy. iTransformer(Liu et al., 2023) addresses scalability by integrating hierarchical attention mechanisms and temporal convolutional networks, capturing both short-term and long-term dependencies more effectively. This lightweight architecture balances accuracy and efficiency, making it suitable for real-world applications. Crossformer (Zhang and Yan, 2022) employs a Dimension-segment-wise structure to embed time series into a two-dimensional vector, effectively capturing dependencies across time and dimensions through its Two-stage-attention structure. The segmentation technique utilized in Crossformer is particularly advantageous for long-term time series forecasting.
@@ 2.2 MLP Models
MLP methods have garnered significant attention in long-term time series forecasting due to their robust performance. Dlinear (Zeng et al., 2023) decomposes time series through the application of moving average kernels and remainder components, yielding two feature variables. These variables are then mapped to the prediction results via linear layers. The Dlinear model can be described as:
$\hat{Y}=W_{1}(MA(X))+W_{2}(R(X))$ (3)
where $MA(X)$ represents the trend component obtained by the moving average kernel, $R(X)$ represents the remainder component, and $W_{1}$, $W_{2}$ are the weights of the linear layers.
TiDE (Das et al., 2023b) introduces a residual structure within the linear layer to encode past time series and covariates, subsequently decoding the encoded time series with future covariates to obtain prediction results. The core operation of TiDE involves projecting features and encoding them through residual blocks, which can be described as:
$\hat{Y}=\text{ResBlock}(\text{ReLU}(W_{f}X+b_{f})+\text{Covariates})$ (4)
where $W_{f}$ and $b_{f}$ are the weights and biases of the projection layer, respectively. The ResBlock operation encapsulates the residual connections used to maintain the integrity of the encoded features while integrating the covariates.
MLP-based methods for long-term time series analysis have demonstrated outstanding predictive accuracy while maintaining low time-space complexity, thus significantly advancing the development of MLP in long-term time series forecasting.
@@ 2.3 CNN Models
The core of time series analysis based on CNN architectures lies in utilizing convolutional kernels to extract temporal and channel features. MICN (Wang et al., 2022) aims to fully exploit the latent information in time series by designing a multi-scale hybrid decomposition module to decompose the complex patterns of input time series. Each module within this design leverages subsampling convolution and equidistant convolution to respectively extract local characteristics and global correlations. TimesNet (Wu et al., 2022) transforms one-dimensional time series into a set of two-dimensional tensors based on multiple periodicities. It further introduces the TimesBlock to adaptively uncover these multi-periodic patterns and effectively extracts intricate temporal dynamics from the transformed tensors through parameter-efficient initializations. These methodologies have attained remarkable results in the domain of long-term time series forecasting.
2.4 SegRNN Model
As the only RNN-based model specifically designed for long-term time series tasks, SegRNN has achieved remarkable performance. SegRNN addresses the challenge of retaining sequence information while minimizing information loss within the recurrent structure through two key methods:
  • SegRNN replaces point-wise iterations with segment-wise iterations. This approach balances the trade-off between sequence information retention and the number of recurrent iterations, leading to improved outcomes.
  • SegRNN replaces recurrent multi-step forecasting with parallel multi-step forecasting. This parallel decoding approach enhances the model’s predictive capability and inference speed.
By combining these two strategies, SegRNN achieves excellent performance within a lightweight model structure, along with superior inference speed.
@@ 3 Method
3.1 Overview
Previous work, such as SegRNN, has achieved remarkable results by employing segmentation and parallel decoding techniques. However, the fixed segmentation approach restricts the model’s capacity and disrupts data continuity. Additionally, segmentation methods still face the challenge of cumulative information processing within the recurrent structure.
In this paper, we propose the MSegRNN method to overcome these limitations through three strategies, as illustrated in Figure 1. We first introduce implicit segmentation, which allows for dense information exchange during the segmentation stage. Subsequently, we incorporate a residual structure to enable partial information bypassing the recurrent structure, thereby reducing information loss in recurrent structure. We also integrate the Mamba architecture to selectively process time series information, further enhancing the model’s capability in handling temporal data.
@@ 3.1 Overview
3.2 Mamba structure
The Mamba architecture, initially proposed by Gu and Dao (2023), has demonstrated outstanding performance in various fields such as large language models and long time series forecasting. By setting model parameters to be content-related, the Mamba architecture achieves powerful data selection capabilities, which are crucial for time series forecasting tasks. Therefore, we utilize the Mamba architecture at the initial stage to preprocess and enhance our time series data. It is important to note that our model does not incorporate the convolutional layer of the Mamba architecture except for weather dataset. This difference is mainly based on performance considerations, as we believe that local information can be effectively captured through segmentation.
@@ 3.3 Implicit Segmentation
The segmentation method employed by SegRNN, which we refer to as explicit segmentation to distinguish it from our segmentation approach, is implemented by truncating the time dimension. For a given time series $X^{(i)}\in\mathbf{R}^{L\times C}$, selecting a segmentation length $w$ results in segmented vectors of size $X_{w}^{(i)}\in\mathbf{R}^{n\times w\times C}$, where $n$ represents the number of segments.
@@
Subsequently, we use linear mapping to transform the dimension $w$ into $d$, obtaining the embedded hidden state of the segmented vectors $\tilde{X}_{w}^{(i)}\in\mathbf{R}^{C\times n\times d}$. This segmentation method retains the sequence information while reducing the iteration count within the recurrent structure. However, this approach has some drawbacks. The simple truncation of data disrupts continuity, and there is no effective exchange of information between different segments, thus limiting the model’s capability.
To address these issues, we propose an implicit segmentation strategy, as illustrated in Figure 2. This strategy initially decomposes the time series through linear mapping and subsequently transforms the information into the embedded hidden state $\tilde{X}_{w}^{(i)}\in\mathbf{R}^{C\times n\times d}$. Notably, after segmentation, each segment includes both the information of that segment and some additional information, ensuring more continuous information processing and thereby enhancing the model’s performance.
Figure 2: Comparsion of SegRNN Segmentation with Implicit Segmentation, we utilize a linear layers, from 1 to $n$, to decompose the time series. Subsequently, a second mapping is applied to compress the dimensions. Note that each segmented vector embedding can encompass information from other segments of the original time series.
Furthermore, our proposed method does not explicitly define the segmentation length. Instead, it achieves data segmentation and dimensionality transformation through dual linear projections. The incorporation of linear fully connected layers facilitates a more continuous and dense transformation of information. It is importanct to emphasize that this implicit segmentation strategy is not confined to a singe model but rather represents a general and straightforward technique for information enhancement.
3.4 Residual Structure
SegRNN aims to minimize the number of iterations in the RNN while preserving sequential information. Extensive experiments show that increasing the number of iterations negatively impacts information retention. SegRNN uses data segmentation to reduce the data’s flow through the RNN structure. Specifically, for the vector obtained through segmentation in SegRNN, $\tilde{X}_{w}^{(i)}\in\mathbf{R}^{C\times n\times d}$, we transform it in the RNN structure with a hidden layer dimension $d$ using the following formula:
@@ 3.4 Residual Structure
where $\mathbf{h}_{t}\in\mathbf{R}^{1\times C\times d}$ represents the hidden state at step $t$, $\tilde{\mathbf{X}}_{w}^{(i)}[:,t,:]$ represents the input vector at step $t$, $\mathbf{W}$ denotes the RNN’s weight parameters, and $f$ is the RNN’s activation function.
Using segmentation techniques, SegRNN reduces the number of iterations from $L$ to $n$. However, the number of iterations still increases with the forecasting time steps, which limits the model’s capability due to the loss of information during these iterations. To address this issue, we incorporate a residual structure into the model. For the segmented vector $\bar{\mathbf{X}}^{(i)}\in\mathbf{R}^{C\times L\times n}$, we apply a linear mapping to $\bar{\mathbf{X}}^{(i)}$, resulting in $\mathbf{h}^{\prime}_{n}$. Finally, we add $\mathbf{h}_{n}$ to the corresponding position in $\mathbf{h}^{\prime}_{n}$ to obtain the final output vector from the encoder.
Introducing the residual structure allows some information to skip the recurrent structure and be directly mapped to the encoder output. This approach minimizes information loss within the recurrent structure, providing a beneficial enhancement for time series data.